The following table shows the number of males and females who are hypertensive and
obese in a research study.
Subject Hypertensive Obese Total
Male 10 42 52
Female 43 5 48
Total 53 47 100
Test whether there is a relationship between hypertension and obesity in the males and females
separately, under the given study.Test at 0.05 level of significance.
1
Expert's answer
2020-04-26T18:52:13-0400
"\chi^2=\frac{(10-27.56)^2}{27.56}+\frac{(42-24.44)^2}{24.44}\\+\frac{(4-25.44)^2}{25.44}+\frac{(5-22.56)^2}{22.56}=55.54"

Critical value:

"\chi^2(1,0.05)= 3.84"

Test statistic is greater than critical value. Thus, we can conclude that there is a relationship between hypertension and obesity in the males and females

separately, under the given study at 0.05 level of significance.